# Vital Insight: Assisting Experts’ Context-Driven Sensemaking of Multi-modal Personal Tracking Data Using Visualization and Human-In-The-Loop LLM Agents

### v2

@@ 5.1.1. Data collection
To generate the dashboard used in the user testing session, we used real-world data from a deployment with participants, where we collected: 1) pre-study surveys, 2) passive mobile and wearable sensing data, and 3) voice assistant check-ins. Participants completed surveys on demographics and regular routines before deployment. We provided the participants with a Garmin smartwatch and installed the study app on their phones. Our app directly syncs with the Garmin device and collects their physical activity, location, call logs, app usage, Wi-Fi/Bluetooth connections, IBI, heart rate, accelerometer, step counts and more. Additionally, participants can initiate a chit-chat conversation with an LLM-powered chatbot on the study provided Amazon Alexa to check in on their day. The system runs in the background, continuously storing and uploading data to our secure server. We used a secure, HIPAA-compliant GPT-4 model that does not store data for retraining. Additionally, we ensured that no identifiable data is transmitted to the model. For example, we only send the de-identified location labels such as ‘home’ to the external API services instead of raw GPS location, and manually removed potentially identifiable Wi-Fi names and user check-ins. The dashboard presented to the experts does not contain any identifiable information. We used tokens to provide temporary access to the system during the user testing. These precautions ensure that neither GPT-4 nor the participants can identify the subject of data collection. The individuals who provided data consented to share it with external researchers through an IRB-approved study.
@@ 5.1.2. Study Design
At the beginning of the session, we provided experts with a brief tutorial on VI-1. We then gave them a link to the system and asked them to share their screen as they completed a series of tasks with the information on the system. As an exploratory session, we used the same day of data for all the user testing sessions for consistency. We designed these tasks to specifically address the four sensemaking scenarios we identified during the initial interviews. For the first task, we asked experts to use all available information on the system and write a short, general summary of the person’s day. We kept the task broad, allowing experts to determine their own version of what the person’s day entailed. Next, experts were asked to dive deeper into specific activities throughout the day in greater detail. We then focused on anomalies, asked experts to identify potential anomalies in the data, and guided them to the corresponding panels to share their thoughts on the results from the system. Finally, we discussed interaction with stakeholders, directing experts to the question section and discussing possible actions they might take based on the information provided. After completing these tasks, we had an open-ended discussion to gather their general thoughts on the system and provided them with a survey. In the survey, we first asked experts to evaluate the different modules of the system. They rated each module based on its importance, trustworthiness, utilization, and clarity.
@@ 5.2.1. Survey results
First, we report the results of the survey. Regarding the impact of different modules on sensemaking. The results indicate that Visualization was clearly the most crucial component, while Hourly Summary was the least, with the other three modules (User Check-in, Daily Summary, and Anomaly) being similarly important. In terms of trust in different modules (scored 1=Distrust to 5=Trust), experts trusted User Check-ins (Mean 4.37, SD=0.47) and Visualization (4.32, SD=0.50) the most, followed by Daily Summary (3.91, SD=0.75), Hourly Summary (3.73, SD=0.69), and Anomaly (3.69, SD=1.03) (Fig. 6). The average trustworthiness score of the system as a whole was 4.09 (SD=0.70). This suggests that the experts generally considered the system to be trustworthy, with the indirect inference components by LLM rated slightly less than the direct data representations. The lower trust in the Anomaly module may also be due to differing definitions of what constitutes an anomaly and vague guidance on how it was determined. Next, we examined the usage and clarity of each data view in the visualization. Among all seven data views in the visualization, experts used an average number of 6.23 (SD=1.01). Heart rate, steps, and activity were used by all experts (13), followed by respiration (12), WiFi (11), battery (10), and phone unlock usage (9). Regarding the clarity of each modality (scored 1 = Very unclear to 5 = Very clear), the average score across all seven modalities was 4.29 (SD=0.37). The clarity ratings from most to least clear were battery (4.77, SD=0.44), WiFi (4.62, SD=0.51), phone unlock usage (4.54, SD=0.52), heart rate (4.23, SD=1.3), respiration (4.08, SD=1.26), steps (4, SD=1.15), and activity (3.77, SD=1.17). These findings show that most data views were generally clear, and experts actively used this information during their sensemaking process.
@@ 5.2.2. General usability and qualitative feedback
Experts loved the visualization aspect of the prototype. For example, P3 specifically mentioned that they appreciated how the steps from both the phone and watch were plotted together. Further, experts found the User Profile module helpful (“The user profile, that’s very powerful(P1)”). Similarly, experts loved the user check-in module and wanted it to be more prominent on the dashboard. Some experts like P1 suggested extracting key information from the user check-in panel and displaying it on the visualization according to the time. For example, P1 suggested it would be helpful ”if the dashboard said, Oh, this is when the little boy sat. This is when the brunch happened.” as an addition to the original paragraph. Additionally, P4 and others suggested displaying timestamps along with some semantics for the conversation with chatbot in the user check-in module, other than only as an event on the visualization.
@@ 6. Final Prototype: Seamless Integration of AI Assistance with Visualization
Furthermore, given the ambiguous definition and low trust in LLM-detected anomalies in Study 2, we now frame “anomalies” as significant Occurrences of interest. Instead of letting the LLM directly detect anomalies without any guidance, we have now introduced a more human-in-the-loop process (Figure.  2). In this approach, we used the experts’ sensemaking process from Study 2 to first identify the important occurrences and used LLMs to provide corresponding explanations. We identified five major occurrence features from Study 2: changes, long durations, discrepancy, gap, and routine activities. For example, we observed that the changes in the location label was a key instance that caught the experts’ attention, prompting them to begin their exploration and examination. Thus, in VI-2 we detect all occurrences where the location label changed and sent the related data around that time period to the LLM to obtain explanations.
@@ 7.1.1. Data collection
We used the same sensing system to collect data. Additionally, to determine whether the sensemaking results from researchers were accurate, participants self-reported fine-grained activities through ecological momentary assessments (EMA) multiple times a day, in addition to the daily check-ins with Alexa. The EMA results typically consisted of a few words describing the current activity, such as ”going out for lunch.” We used EMAs solely as the ground truth for momentary user behavior to determine the accuracy of the if the sensemaking results from researchers. We did not provide the EMAs to the LLMs or the experts.
@@ 7.1.2. Study Design
We conducted an A/B test to investigate whether LLM assistance in VI-2 is helpful for researchers in performing sensemaking tasks. Condition A used VI-2 with full LLM assistance, while Condition B used VI-2 without LLM assistance. In Condition B, the experts did not have access to the Day-in-a-Glance and the Occurrences explanation. During the experiment, each participant interacted with data from four different days, with two days for each condition, and answered a series of questions. The order of the tasks (A/B conditions and order of days) was randomized to ensure counterbalance and avoid carry-over effect. For each day, researchers first answered three multiple-choice questions related to the person’s day using all the information provided in the system. For example, a sample question could be ”What is the person doing at 6:30 pm?”, with several options like ”going out for lunch”, ”hanging out with friends”, etc. After answering each question, experts rated their experience by completing the Single Usability Metrics survey (SUM) across the dimensions of ease, time, and satisfaction (sauro2005method, ), along with two Likert-scale questions to rate the AI component (only in Condition A, where LLM is enabled). Next, participants wrote a short summary of the person’s day within 3 minutes and completed the same surveys they filled out after answering the previous selection questions. After each task, participants discussed their thought processes as qualitative feedback for the research team. The questions module was integrated directly into the interface during the user testing  4. After the experiment, the participants filled out two System Usability Scale (SUS) to rate the overall usability of LLM enabled and disabled prototypes (brooke1996sus, ), and answered several questions related to the helpfulness of AI assistance. At the end of the study, we asked participants several questions aimed at identifying potential actions based on the four days of data. For example, we asked what insights could be generated about the person, and inquired about the types of health interventions they might design to improve the person’s physical activity levels. Since the correctness of the answers could not be objectively measured during this user testing, we kept this as an open-ended discussion with the experts to gather qualitative insights.
@@ 7.2.1. Descriptive Results
First, we discuss the descriptive results from the study. For correctness in answering the questions, we summed the scores for each participant to calculate a final score. For answers that were completely correct, we gave a score of 1; for those that were partially correct (e.g., missing one option in a multi-option question), we gave a score of 0.5; and for answers with any incorrect option, we gave a score of 0. Each participant answered 3 questions for each of the 4 days thus the final score ranged from 0 to 12. The final scores of 20 experts ranged from 1.5 to 11, with outliers (Median = 8.5, SD = 2.34). We believe there is no ceiling effect for the questions, and the study and question design are appropriate to capture some differences between the two conditions. For each question where LLM was enabled, we asked participants to rate the trustworthiness and helpfulness of the AI component (scored 1=Distrust/Unhelpful to 5=Trust/Helpful),. The mean trustworthiness and helpfulness scores were 3.94 (SD = 0.80) and 4 (SD = 0.91), respectively, indicating a generally good experience with the LLM (4 = Somewhat Trustworthy/Helpful).
@@ 7.2.2. Results of A/B test
The distribution of SUS scores with LLM followed a normal distribution (mean = 73.42, SD = 17.92), while the distribution without LLM was negatively skewed (mean = 70.53, median = 75, SD = 19.59). The mean SUS scores were considered ”Good” and ”Acceptable” for the LLM-enabled condition, and ”OK” and ”Marginal High” for the LLM-disabled condition (bangor2009determining, ; bangor2008empirical, ). Using a Wilcoxon Signed-Rank Test, we found that having AI-assistance led significantly higher SUS scores than without AI-assistance($p<0.05$), with a medium effect (effect size = 0.44).
@@ 7.2.5. User Interaction: Experts’ sensemaking process
In Study 3, we posed specific questions for experts to answer. This approach revealed new insights to the sensemaking process that is complementary to the more open-ended summary-writing scenario in Study 2. Some experts immediately began identifying related data based on the question. The first thing they tried to identify was time, which helped them narrow down the data set to focus on. Some questions also prompted experts to focus on a specific modality. For instance, when asked what the person was doing during a long phone usage period, many experts went directly to the phone usage data and only turned to other modalities as necessary. In such cases, LLM occurrences became an additional modality, on par with other sensing modalities, assisting experts to make sense of what happened. Interestingly, despite the specific questions, some experts still started by reviewing the entire day’s data to first get an overview. This process is similar to what we observed in Study 2. In this scenario, both LLM occurrences and the day-in-a-glance feature were the first things some experts examined, especially when they were familiar with LLM and ”knew what to expect.”

### v3

@@ 5.1.1. Data collection
To generate the dashboard used in the user testing session, we used real-world data from a deployment with participants, where we collected: 1) pre-study surveys, 2) passive mobile and wearable sensing data, and 3) voice assistant check-ins. Participants completed surveys on demographics and regular routines before deployment. We provided the participants with a Garmin smartwatch and installed the study app on their phones. Our app directly syncs with the Garmin device and collects their physical activity, location, call logs, app usage, Wi-Fi/Bluetooth connections, IBI, heart rate, accelerometer, step counts and more. Additionally, participants can initiate a chit-chat conversation with an LLM-powered chatbot on the study provided Amazon Alexa to check in on their day. The system has been deployed in prior studies (campbell2023patient, ; mishra2021detecting, ; kunzler2019exploring, ; mishra2022towards, ; aip_2024, ), and runs in the background, continuously storing and uploading data to our secure server. We used a secure, HIPAA-compliant GPT-4 model that does not store data for retraining. Additionally, we ensured that no identifiable data is transmitted to the model. For example, we only send the de-identified location labels such as ‘home’ to the external API services instead of raw GPS location, and manually removed potentially identifiable Wi-Fi names and user check-ins. The dashboard presented to the experts does not contain any identifiable information. We used tokens to provide temporary access to the system during the user testing. These precautions ensure that neither GPT-4 nor the participants can identify the subject of data collection. The individuals who provided data consented to share it with external researchers through an IRB-approved study.
@@ 5.1.2. Study Design
At the beginning of the session, we provided experts with a brief tutorial on VI-1. We then gave them a link to the system and asked them to share their screen as they completed a series of tasks with the information on the system. As an exploratory session, we used the same day of data for all the user testing sessions for consistency. We designed these tasks to specifically address the four sensemaking scenarios we identified during the initial interviews. For the first task, we asked experts to use all available information on the system and write a short, general summary of the person’s day. We kept the task broad, allowing experts to determine their own version of what the person’s day entailed. Next, experts were asked to dive deeper into specific activities throughout the day in greater detail. We then focused on anomalies, asked experts to identify potential anomalies in the data, and guided them to the corresponding panels to share their thoughts on the results from the system. Finally, we discussed interaction with stakeholders, directing experts to the question section and discussing possible actions they might take based on the information provided. After completing these tasks, we had an open-ended discussion to gather their general thoughts on the system and provided them with a survey. In the survey, we first asked experts to evaluate the different modules of the system. They rated each module based on its importance, trustworthiness, utilization, and clarity, common metrics in evaluation visualization based dashboard (10360907, ; almasi2023usability, ).
@@ 5.2.1. Survey results
First, we report the results of the survey. Regarding the impact of different modules on sensemaking. The results indicate that Visualization was clearly the most crucial component, while Hourly Summary was the least, with the other three modules (User Check-in, Daily Summary, and Anomaly) being similarly important. In terms of trust in different modules (scored 1=Distrust to 5=Trust), experts trusted User Check-ins (Mean 4.37, SD=0.47) and Visualization (4.32, SD=0.50) the most, followed by Daily Summary (3.91, SD=0.75), Hourly Summary (3.73, SD=0.69), and Anomaly (3.69, SD=1.03) (Fig. 7). The average trustworthiness score of the system as a whole was 4.09 (SD=0.70). This suggests that the experts generally considered the system to be trustworthy, with the indirect inference components by LLM rated slightly less than the direct data representations. The lower trust in the Anomaly module may also be due to differing definitions of what constitutes an anomaly and vague guidance on how it was determined. Next, we examined the usage and clarity of each data view in the visualization. Among all seven data views in the visualization, experts used an average number of 6.23 (SD=1.01). Heart rate, steps, and activity were used by all experts (13), followed by respiration (12), WiFi (11), battery (10), and phone unlock usage (9). Regarding the clarity of each modality (scored 1 = Very unclear to 5 = Very clear), the average score across all seven modalities was 4.29 (SD=0.37). The clarity ratings from most to least clear were battery (4.77, SD=0.44), WiFi (4.62, SD=0.51), phone unlock usage (4.54, SD=0.52), heart rate (4.23, SD=1.3), respiration (4.08, SD=1.26), steps (4, SD=1.15), and activity (3.77, SD=1.17). These findings show that most data views were generally clear, and experts actively used this information during their sensemaking process, which aligns with other studies on visualization systems that also consider a similar score to be a good rating (10.1145/3613904.3642237, ; 10.1145/3613904.3643022, ; zhu2007measuring, ; north2006toward, ).
@@ 5.2.2. General usability and qualitative feedback
Experts loved the visualization aspect of the prototype. For example, P3 specifically mentioned that they appreciated how the steps from both the phone and watch were plotted together. Further, experts found the User Profile module helpful (“The user profile, that’s very powerful(P1)”). Similarly, experts loved the user check-in module and wanted it to be more prominent on the dashboard. Some experts like P1 suggested extracting key information from the user check-in panel and displaying it on the visualization according to the time. For example, P1 suggested it would be helpful “if the dashboard said, Oh, this is when the little boy sat. This is when the brunch happened.” as an addition to the original paragraph. Additionally, P4 and others suggested displaying timestamps along with some semantics for the conversation with chatbot in the user check-in module, other than only as an event on the visualization.
@@ 5.2.3. User interaction: Experts’ sensemaking process
5.3. Follow-up evaluation on LLM anomaly detection
Through the interviews, we discovered that the LLM’s anomaly detection on a single modality did not align with the experts’ expectations. As a result, we conducted an exploratory analysis to evaluate LLM anomaly detection. We ran the LLM anomaly detection on heart rate and respiration data for a week (11/18/2024 - 11/24/2024), performing 10 tests for each day using the same prompts. We observed significant instability in the detection on several days (Figure.  5). Ideally, all the different colors of the horizontal lines should align for the same day. In our evaluation, although there is some consistency on several days (e.g., heart rate anomaly on 11/20), many outputs are scattered across the 24-hour period. Additionally, there are two significant long heart rate anomaly durations in the initial LLM outputs on 11/23 and 11/19 that are inaccurate. As a result, the LLM anomaly detection results not only misalign with the experts’ mental model but are also highly inconsistent and unstable. This highlights the need for a more expert-instructed anomaly detection process to ensure greater stability.
Figure 5. Time ranges of LLM detected heart rate and respiration anomalies from 11/18/2024 to 11/24/2024. The horizonal lines represents the start and end time of the anomalies, and the different colors represent 10 LLM test runs.
[…]
@@ 6. Final Prototype: Seamless Integration of AI Assistance with Visualization
Furthermore, given the ambiguous definition and low trust in LLM-detected anomalies in Study 2, we now frame “anomalies” as significant Occurrences of interest. Instead of letting the LLM directly detect anomalies without any guidance, we have now introduced a more human-in-the-loop process to mitigate potential inconsistencies and inaccurate results in LLM anomaly detection (Figure.  2). In this approach, we used the experts’ sensemaking process from Study 2 to first identify the important occurrences, which are typically preprocessed features. These features, along with the semantic descriptions of all modalities of sensing data around that time (as described in Section 4.3), are then sent to the LLMs for corresponding explanations. We identified five major occurrence features from Study 2: changes, long durations, discrepancy, gap, and routine activities. For example, we observed that the changes in the location label was a key instance that caught the experts’ attention, prompting them to begin their exploration and examination. Thus, in VI-2 we detect all occurrences where the location label changed and sent the related data around that time period to the LLM to obtain explanations.
@@ 7.1.1. Data collection
We used the same sensing system to collect data. Additionally, to determine whether the sensemaking results from researchers were accurate, participants self-reported fine-grained activities through ecological momentary assessments (EMA) multiple times a day, in addition to the daily check-ins with Alexa. The EMA results typically consisted of a few words describing the current activity, such as “going out for lunch.” We used EMAs solely as the ground truth for momentary user behavior to determine the accuracy of the if the sensemaking results from researchers. We did not provide the EMAs to the LLMs or the experts.
@@ 7.1.2. Study Design
We conducted an A/B test to investigate whether LLM assistance in VI-2 is helpful for researchers in performing sensemaking tasks. Condition A used VI-2 with full LLM assistance, while Condition B used VI-2 without LLM assistance. In Condition B, the experts did not have access to the Day-in-a-Glance and the Occurrences explanation. During the experiment, each participant interacted with data from four different days, with two days for each condition, and answered a series of questions. The order of the tasks (A/B conditions and order of days) was randomized to ensure counterbalance and avoid carry-over effect. For each day, researchers first answered three multiple-choice questions related to the person’s day using all the information provided in the system. For example, a sample question could be “What is the person doing at 6:30 pm?”, with several options like “going out for lunch”, “hanging out with friends”, etc. After answering each question, experts rated their experience by completing the Single Usability Metrics survey (SUM) across the dimensions of ease, time, and satisfaction (sauro2005method, ), along with two Likert-scale questions to rate the AI component (only in Condition A, where LLM is enabled). Next, participants wrote a short summary of the person’s day within 3 minutes and completed the same surveys they filled out after answering the previous selection questions. After each task, participants discussed their thought processes as qualitative feedback for the research team. The questions module was integrated directly into the interface during the user testing  4. After the experiment, the participants filled out two System Usability Scale (SUS) to rate the overall usability of LLM enabled and disabled prototypes (brooke1996sus, ), and answered several questions related to the helpfulness of AI assistance. At the end of the study, we asked participants several questions aimed at identifying potential actions based on the four days of data. For example, we asked what insights could be generated about the person, and inquired about the types of health interventions they might design to improve the person’s physical activity levels. Since the correctness of the answers could not be objectively measured during this user testing, we kept this as an open-ended discussion with the experts to gather qualitative insights.
@@
7.2. Preliminary evaluation on LLM summary
Although not the primary focus of our research, we conducted a preliminary evaluation of the LLM summary to assess its consistency, accuracy, and processing overhead to gain an overview of the system performance. We ran the LLM backend 10 times over data from four days and compared the LLM-generated summary with summaries written by multiple experts.
For consistency, we calculated the similarity between summaries generated for the same day of data using the Term Frequency-Inverse Document Frequency (TF-IDF) method, combined with Cosine Similarity to avoid over-sensitivity to word frequency. We calculated the similarity matrix between each pair of LLM summaries (45 pairs for 10 summaries per day) and summaries written by different experts. The mean similarity of same-day summaries generated by LLM (Mean = 0.52, SD = 0.15) is significantly higher than those written by experts (Mean = 0.23, SD = 0.11), which aligns with the first author’s observation. However, we want to emphasize that the summary of daily activities is highly subjective, and traditional NLP metrics like similarity might not perfectly capture the nuances of this type of task.
Regarding accuracy, the first author manually reviewed each fact in the LLM and expert summaries, cross-referencing with passive tracking and EMA data. For example, the statement “They appear to have gone to bed around 12:30 am and woke up around 9:15 am” contains two facts: “gone to bed around 12:30 am” and “woke up around 9:15 am.” Each fact was categorized as ‘correct’ if it aligned with self-report or sensing data, ‘wrong’ if it contradicted the data, and ‘unclear’ if the fact could not be conclusively determined from the available data. For the LLM summary, there were 5 wrong facts (2.13%), 215 correct facts (91.49%), and 15 unclear facts (6.38%) out of 235 facts. For the expert summary, there were 15 wrong facts (11.03%), 114 correct facts (83.82%), and 7 unclear facts (5.15%) out of 136 facts. LLM sometimes makes mistakes such as overstating details (e.g., “Maintained a WiFi connection throughout the day” when the WiFi was disconnected for a brief period), mismatches and hallucinations in time (e.g., “Casual Shopping: User engaged in casual shopping activities, enhancing the evening’s enjoyment.” when shopping actually occurred in the morning). It also tends to draw overly aggressive connections between data (e.g., “showing varied heart rate and respiration patterns while connected to WiFi” when no clear relationship exists), and makes ‘reasonable’ but slightly overstated and unnecessary guesses (e.g., “The user frequently interacted with their phone throughout the day, suggesting engagement with social media, messaging, or other applications,” despite no tracking of social media usage). Additionally, we observed an instance where the LLM had issues with numbers when summarizing heart rate data due to the loss of detailed raw data under the multi-level processing structure. For experts, many mistakes stem from missing data (e.g., ignoring details from user check-ins or profiles) or misinterpreting complex sensemaking between multiple modalities (e.g., wrongly identifying sleep when no steps are detected, but ignoring active phone usage data).
While the LLM summary had a higher accuracy rate, we found that expert summaries were generally more detailed, often including specific times, while LLM summaries tended to be more generalized about daily routines. For example, experts might summarize with “They came back home around 7pm,” while LLM might write “Came back home later in the evening.” We calculated the ratio of the number of facts included in the summary to the total number of tokens. The mean ratio was 5% (SD = 1%) for LLM-generated summaries and 9% (SD = 2%) for expert-generated summaries. We will discuss more on the qualitative feedback from experts in the next section.
Finally, we calculated the overhead for LLM usage. On average, there were 262182.25 input tokens and 1241.83 output tokens processed per day. For the GPT-4o-mini model used in Vital Insight, the API cost was \$0.0397 per day, approximately \$14.50 per year for a longitudinal deployment. Although the price is reasonable, it could increase if the system needs to run multiple times to avoid inconsistencies.
@@ 7.3.1. Descriptive Results
First, we discuss the descriptive results from the study. For correctness in answering the questions, we summed the scores for each participant to calculate a final score. For answers that were completely correct, we gave a score of 1; for those that were partially correct (e.g., missing one option in a multi-option question), we gave a score of 0.5; and for answers with any incorrect option, we gave a score of 0. Each participant answered 3 questions for each of the 4 days thus the final score ranged from 0 to 12. The final scores of 20 experts ranged from 1.5 to 11, with outliers (Median = 8.5, SD = 2.34). We believe there is no ceiling effect for the questions, and the study and question design are appropriate to capture some differences between the two conditions. For each question where LLM was enabled, we asked participants to rate the trustworthiness and helpfulness of the AI component (scored 1=Distrust/Unhelpful to 5=Trust/Helpful),. The mean trustworthiness and helpfulness scores were 3.94 (SD = 0.80) and 4 (SD = 0.91), respectively, indicating a generally good experience with the LLM (4 = Somewhat Trustworthy/Helpful), which aligns with other studies on AI-assisted systems that also consider a similar score to be a good rating  (yang2024talk2care, ; 10296017, ).
@@ 7.3.2. Results of A/B test
The distribution of SUS scores with LLM followed a normal distribution (mean = 73.42, SD = 17.92), while the distribution without LLM was negatively skewed (mean = 70.53, median = 75, SD = 19.59). The mean SUS scores were considered “Good” and “Acceptable” for the LLM-enabled condition, and “OK” and “Marginal High” for the LLM-disabled condition (bangor2009determining, ; bangor2008empirical, ). Using a Wilcoxon Signed-Rank Test, we found that having AI-assistance led significantly higher SUS scores than without AI-assistance($p<0.05$), with a medium effect (effect size = 0.44).
@@ 7.3.5. User Interaction: Experts’ sensemaking process
In Study 3, we posed specific questions for experts to answer. This approach revealed new insights to the sensemaking process that is complementary to the more open-ended summary-writing scenario in Study 2. Some experts immediately began identifying related data based on the question. The first thing they tried to identify was time, which helped them narrow down the data set to focus on. Some questions also prompted experts to focus on a specific modality. For instance, when asked what the person was doing during a long phone usage period, many experts went directly to the phone usage data and only turned to other modalities as necessary. In such cases, LLM occurrences became an additional modality, on par with other sensing modalities, assisting experts to make sense of what happened. Interestingly, despite the specific questions, some experts still started by reviewing the entire day’s data to first get an overview. This process is similar to what we observed in Study 2. In this scenario, both LLM occurrences and the day-in-a-glance feature were the first things some experts examined, especially when they were familiar with LLM and “knew what to expect.”